Reports 1-1 of 1 Clear search Modify search
MIF (General)
keiko.kokeyama - 17:57 Tuesday 12 April 2016 (1506) Print this report
MICH sensor changed/ Actuator cal lines

During the little commissioning brake today, following changes were applied for MICH configuration. They are valid from 4/12 16:10:00 JST.

  • Three cal lines were injected to EX, EY and BS to monitor the actuator efficiency (111, 113 and 115Hz for EX, EY and BS respectively). OSC and demod names are such as K1:LSC-ACT_OSC1_CLKGAIN and K1:LSC-ACT_DEMOD1_I_OUT and so on. ACT means actuator. According to this change, UGF SERVO (main) cal lines were renamed such as K1:LSC-CAL_DEMOD2_I_OUT (LKIN was removed from the channel name) so that channel names have coherence.
  • MICH sensor was changed from I to Q. Phase is adjusted to maximize Q signal. Optical gain (which is now not used for the calibration) is same as before (klog1455)
  • With Q sensor, UGF (without UGF servo) is confirmed as 100Hz. UGF servo is on to compensate the optical gain fluctuation.
Comments to this report:
keiko.kokeyama - 17:38 Monday 18 April 2016 (1569) Print this report

Here are the plots for error signals before (OUT) and after (ERR) the demod phase adjustment. 

Images attached to this comment
keiko.kokeyama - 15:09 Tuesday 19 April 2016 (1586) Print this report

OLT plots after the sensor was changed to Q.

Non-image files attached to this comment
Search Help